# CAPTURe: Evaluating Spatial Reasoning in Vision Language Models via Occluded Object Counting

### v1

@@ Abstract
Recognizing and reasoning about occluded (partially or fully hidden) objects is vital to understanding visual scenes, as occlusions frequently occur in real-world environments and act as obstacles for spatial comprehension. To test models’ ability to reason about multiple occluded objects, we introduce a novel task, Counting Amodally for Patterns Through Unseen REgions (CAPTURe), which requires a model to count objects arranged in a pattern by inferring how the pattern continues behind an occluder (an object which blocks parts of the scene). CAPTURe requires both recognizing visual patterns and reasoning, making it a useful testbed for evaluating vision-language models (VLMs) on whether they understand occluded patterns and possess spatial understanding skills. By requiring models to reason about occluded objects, CAPTURe also tests VLMs’ ability to form world models that would allow them to fill in missing information. CAPTURe consists of two parts: (1) CAPTURe, with manually filtered images of real objects in patterns and (2) CAPTURe, a controlled diagnostic with generated patterned images. We evaluate four strong VLMs (GPT-4o, Intern-VL2, Molmo, and Qwen2-VL) on CAPTURe, finding that models struggle to count on both occluded and unoccluded patterns. Crucially, we find that models perform worse with occlusion, suggesting that VLMs are also deficient in inferring unseen spatial relationships: even the strongest VLMs like GPT-4o fail to count with occlusion. In contrast, we find that humans achieve very little error on CAPTURe. We also find that providing auxiliary information of occluded object locations increases performance, underscoring that the model error comes both from an inability to handle occlusion as well as difficulty counting in images.111Code and data: https://github.com/atinpothiraj/CAPTURe
@@ 1 Introduction
Inferring what lies behind different objects in occluded scenes is crucial for human perception, as it allows us to maintain a coherent understanding of our environment even when parts are hidden. The human visual system accomplishes this by integrating past experiences, context, and sensory inputs to reconstruct incomplete scenes [30, 27, 44, 19]. Meanwhile, recent advancements in vision-language models (VLMs) – especially in terms of visual and spatial reasoning – raise the question of whether these systems can perform similar inferential tasks. One way of measuring such capabilities is through amodal completion – the task of inferring the invisible parts of partially occluded objects; here, vision-only models are typically evaluated via dense prediction tasks like object segmentation and image inpainting [5]. However, this format is not well-suited for assessing VLMs, whose outputs consist of text tokens rather than pixel-level predictions. This raises a critical question: how can we quantify the ability of VLMs to form spatial world modeling [17] in the presence of occlusion.
To address this, we introduce CAPTURe, Counting Amodally for Patterns Through Unseen REgions, a novel benchmark that tests a VLM’s world modeling and spatial reasoning abilities through the task of amodal counting, where models are prompted to count occluded objects by amodally completing a pattern. CAPTURe focuses on counting as it provides an objective and easy-to-verify output by comparing predicted counts with ground truth values. Moreover, patterned objects appear in a variety of real-world domains, especially in man-made environments like parking lots, cities, and warehouses, where counting objects is often required. Fig. 1 illustrates the CAPTURe task. We show a VLM an image where objects are placed in a regular pattern (e.g., a 4x4 grid) with some objects occluded, and ask the model to count the total number of objects in the image assuming that the pattern continues behind the occlusion. The task requires handling occlusion, pattern recognition, and counting skills that exist in humans from a fairly young age [30, 27, 44], thus humans can easily answer such questions – indeed, we find that people can complete CAPTURe tasks with almost no error.
@@
By combining vision encoders with large language models (LLMs), VLMs have the potential to reason in a zero-shot way about visual inputs. To put this ability to the test and measure VLMs’ ability to reason about missing visual information, we evaluate four strong recent VLMs (GPT-4o, InternVL2, Molmo, and Qwen2VL) on CAPTURe. Our experiment results (Sec. 4) show that models generally struggle with the multiple aspects of task, with high error rates on both CAPTURe and CAPTURe for occluded and unoccluded images. In contrast, we find that humans can perform the task easily: whereas model performance deteriorates as more objects in images are occluded, humans complete the task almost perfectly. We also compare VLMs to a vision-only model trained to count visible objects; while this model generally outperforms VLMs, its error is directly tied to the number of occluded objects – the more objects are occluded, the higher its error will be.
By objectively measuring VLMs’ spatial reasoning capabilities under occlusion, CAPTURe highlights an unexpected weakness in VLMs. We analyze this weakness by providing the model with additional clues and information. Specifically, we test to what degree the VLMs’ failure stems from an inability to integrate visual information by providing it with a text-based representation of the visible objects in the image in the form of object coordinates; here, VLMs perform substantially better, indicating that their poor performance on CAPTURe stems partly from an inability to count objects in images, rather than an inability to count more generally. This aligns with previous work, which similarly finds that VLMs struggle to count in images [33, 22, 41]. We also test the degree to which VLM errors stem from an inability to form a world model by providing it with auxiliary information (the coordinates of the occluded objects in text, or inpainting of occluded regions). We find that VLMs perform substantially better with this auxiliary information, suggesting that VLMs are partly limited by their inability to imagine the missing visual information. Addressing these gaps is critical for VLMs to function effectively in real-world scenarios, where visual reasoning often involves occlusions – whether counting stadium seats, components on production lines, or buildings in neighborhoods. We hope that our work will foster future research on improving the world modeling capabilities of VLMs.
@@ 2.1 Task Overview
CAPTURe tests VLMs in their ability to reason over occluded scenarios, recognize patterns, and count visible and occluded objects. VLMs already achieve high accuracy in classifying single, occluded objects [20]. Thus, we also argue that VLMs have the potential to perform well on CAPTURe’s challenging task because their proficiency in handling occlusion ought to enable them to recognize occluded objects and reason accordingly. All images in CAPTURe contain a pattern. This makes the task solvable for models and people – if the objects were not placed in a pattern, it would be unreasonable to expect models to infer the position of the occluded objects. For example, given an image of a random pile of coins with a region occluded, it is not easy to infer whether the occluded region simply contains no coins or contains roughly the same amount as the rest of the pile. For this task, the patterns considered are all regular and fairly small patterns, e.g. grids, circles, triangles, and other regular shapes – see Fig. 2 for further examples. The last step of CAPTURe is counting, asking the model to provide an objectively measurable output. In addition to VLMs, we also test CountGD [3], a state-of-the-art object detection-based counting method, finding that it fails to account for the occluded scenario, as its training entails solely predicting the visible, unoccluded objects in the image.
@@
where represents the actual values, represents the predicted values, and is the number of observations. sMAPE is capped at 100%, providing a fixed range. This makes sMAPE ideal for challenging tasks like ours, as we can penalize responses that fail to produce an answer with a max error of 100%. For a justification of sMAPE over other metrics, see Sec. A.1.
@@ 2.2 Dataset
We introduce a set of real images with patterns to test amodal counting in naturalistic settings. The original images and annotations come from the FSC-147 dataset [37], a diverse counting dataset with manual annotations for the number of target objects and all object bounding boxes in each image. FSC-147 contains a diverse array of objects, with 6146 real-world images across 147 object categories. We filter FSC-147 for images that contain identifiable and regular patterns of objects and manually overlay a black box to occlude some objects, resulting in 924 images. Filtering is first performed by GPT-4o and then manually verified; we also manually verify that determining objects despite the occlusion is feasible. For each example, we maintain both an occluded version and an unoccluded version. Additional details on how we create CAPTURe can be found in Appendix B.
@@ 2.3 Statistics and Examples
Fig. 2 shows examples from CAPTURe and CAPTURe paired with their corresponding answers from GPT-4o and their ground truth answers. These examples show the range of objects and patterns present in the dataset, and highlight the feasibility of the task for a human. Tab. 1 reports summary statistics for CAPTURe, including the number of images and object types, as well as the mean number of occluded and total objects in both splits of CAPTURe. The number of objects in CAPTURe is shown in Fig. 3, where most images have between 0 and 30 objects. On CAPTURe, the maximum number of objects is 15, and CAPTURe images generally have 1-6 occluded objects (shown in Fig. 4, as further occlusion could make the count unresolvable).
@@ 3.1 Models
We experiment with GPT-4o [28], Intern-VL2-Llama3-8B [9, 10], and Qwen2-VL-7B [40] for their high scores on other VLM tasks [29]. We add Molmo 7B-D [13], because of its ability to “point and count,” giving it a potential advantage on CAPTURe. Specifically, Molmo is trained on millions of examples that directly ground text to 2D coordinates (or “points”) in images. This allows Molmo to directly point to image coordinates, also enabling it to count more easily by pointing to several objects. All the VLMs feature a different language backbone and vision encoder to provide broad coverage of model architectures. To evaluate models, we provide the model with the name of the specific object to be counted and the explicit instruction to count fully visible objects and objects behind the occluding box (in the occluded images). For each model, we test ten prompts on a validation set of 100 images, selecting the best prompt for each model in each dataset section (CAPTURe/CAPTURe) and for each environment (occluded/unoccluded). We provide the selected prompts in Appendix D.
@@ 3.2 Answer Generation and Extraction
Given the complex nature of CAPTURe, we allow models to generate open-ended responses and then subsequently extract answers. Further details (including max number of tokens) can be found in Sec. A.2.
Answer extraction. Empirically, we found that constraining the output to a specific format for ease of analysis negatively impacted benchmark performance. Therefore, we instead prompt models to generate freely and extract the final output number using a separate answer extractor based on Llama 3.1 8B [1]. This answer extractor takes as input the output from the model and is prompted to extract a single number representing the final answer. The answer extractor also identifies if an output failed to converge on a singular number answer and assigns a label to these examples. We mark such incomplete/incoherent model generations as ‘skipped’ questions and when calculating the error later, these responses are assigned the worst possible sMAPE score (100%). The answer extractor outputs were manually verified on 1000 outputs, where the extractor was found to have 100% accuracy.
Human and object detection baselines. We also report the performance of humans and a recent counting model (CountGD [3]) as baselines to establish a point of reference for model performance. To confirm that humans can perform the CAPTURe task, we provided 100 randomly-selected occluded examples each from the CAPTURe and CAPTURe subsets to 3 undergraduate students with no prior knowledge of the task.
@@ 4.1 Main Results on CAPTURe
GPT-4o | 14.09 | 8.40 |
InternVL2 | 29.74 | 24.98 |
Molmo | 33.68 | 17.24 |
Qwen2VL | 29.80 | 9.39 |
Avg. of 4 VLMs | 26.83 | 15.00 |
Models consistently struggle with counting and perform worse on occluded images.
@@
We run the VLMs on the occluded and unoccluded versions of CAPTURe to discern whether occlusion significantly impacts model performance. Tab. 2 shows that all models struggle with counting generally, performing poorly on both splits. Moreover, we see that every model performs better on the unoccluded images. On average, the models perform 6.28% worse in CAPTURe occluded images and 4.85% worse in CAPTURe occluded images (in terms of absolute sMAPE), indicating increased difficulty from a standard counting task. The best model for both splits, GPT-4o, has an error rate of 14.75% on CAPTURe and a lower error rate of 5.90 on synthetic. Across both the real and synthetic split, GPT-4o’s error increases with occlusion, by 1.41% on the real data and 3.81% on the synthetic split. Interestingly, despite its fine-tuning on counting tasks, Molmo exhibits a sizable error rate of 32.5% on CAPTURe occluded images. The high error rates of VLMs indicate limited capabilities in visual understanding under occlusions, pattern recognition, and counting. We further analyze the source of these errors with oracle experiments in Sec. 4.3.
Humans complete the task with almost no error.
@@
Tab. 3, evaluated on a 100-example subset of each split, confirms that humans complete the task with ease despite occlusion, with an sMAPE of 3.79% on CAPTURe and 0.92% on CAPTURe. On the same subset of examples, models performed 7 times worse on CAPTURe and 16 times worse on CAPTURe than humans, underscoring the gap between VLM and humans in this task.
Object detection-based baseline outperforms VLMs.
@@
We attempt the task with a strong object detection-based model to highlight that a standard counting approach will experience a greater loss going from unoccluded to occluded environments, as it cannot capture any occluded objects, i.e. cannot reason. We choose CountGD [3], the top solution for unoccluded counting on FSC-147, on which it was trained. Because we draw our images from FSC-147’s train and test sets, and CountGD trains on FSC-147, we only evaluate CountGD on the subset of our data sourced from the FSC-147 test set, consisting of 149 images.
@@
We find that CountGD deteriorates by 7.19% on occluded images, increasing from 3.15% sMAPE to 10.34% as observed in Fig. 5. As expected, CountGD outperforms all of the VLMs on the unoccluded split as it is trained for counting on FSC-147. CountGD also outperforms the VLMs on the occluded split, reinforcing that only counting the visible objects is a hard-to-beat baseline. However, the drop in performance with occlusion is greater than the average VLM’s drop, highlighting a disadvantage of non-reasoning solutions on CAPTURe: their error is necessarily tied directly to the number of occluded objects and they cannot address the task on their own, whereas a VLM might be able to infer missing objects via reasoning.
Hybrid VLM counting systems improve performance.
@@
Finding that CountGD is far better at counting visible objects than VLMs, we leverage the advantage that CountGD has by feeding its visible object count information to the VLMs as part of the prompt. As expected, Fig. 5 illustrates that there is a considerable decrease in error when CountGD and the VLMs are combined into a hybrid system. However, overall the hybrid system still performs worse than CountGD alone, indicating VLMs are still subpar even at counting just occluded objects (this is further reinforced by Sec. C.2).

### v2

@@ Abstract
Recognizing and reasoning about occluded (partially or fully hidden) objects is vital to understanding visual scenes, as occlusions frequently occur in real-world environments and act as obstacles for spatial comprehension. To test models’ ability to reason about multiple occluded objects, we introduce a novel task, Counting Amodally for Patterns Through Unseen REgions (CAPTURe), which requires a model to count objects arranged in a pattern by inferring how the pattern continues behind an occluder (an object which blocks parts of the scene). CAPTURe requires both recognizing visual patterns and reasoning, making it a useful testbed for evaluating vision-language models (VLMs) on whether they understand occluded patterns and possess spatial understanding skills. By requiring models to reason about occluded objects, CAPTURe also tests VLMs’ ability to form world models that would allow them to fill in missing information. CAPTURe consists of two parts: (1) CAPTURe, with manually filtered images of real objects in patterns and (2) CAPTURe, a controlled diagnostic with generated patterned images. We evaluate four strong VLMs (GPT-4o, Intern-VL2, Molmo, and Qwen2-VL) on CAPTURe, finding that models struggle to count on both occluded and unoccluded patterns. Crucially, we find that models perform worse with occlusion, suggesting that VLMs are also deficient in inferring unseen spatial relationships: even the strongest VLMs like GPT-4o fail to count with occlusion. In contrast, we find that humans achieve very little error on CAPTURe. We also find that providing auxiliary information of occluded object locations increases performance, underscoring that the model error comes both from an inability to handle occlusion as well as difficulty in counting in images.111Code and data: https://github.com/atinpothiraj/CAPTURe
@@ 1 Introduction
Inferring what lies behind different objects in occluded scenes is crucial for human perception, as it allows us to maintain a coherent understanding of our environment even when parts are hidden. The human visual system accomplishes this by integrating past experiences, context, and sensory inputs to reconstruct incomplete scenes [30, 27, 45, 19]. Meanwhile, recent advancements in vision-language models (VLMs) – especially in terms of visual and spatial reasoning – raise the question of whether these systems can perform similar inferential tasks. One way of measuring such capabilities is through amodal completion – the task of inferring the invisible parts of partially occluded objects; here, vision-only models are typically evaluated via dense prediction tasks like object segmentation and image inpainting [5]. However, this format is not well-suited for assessing VLMs, whose outputs consist of text tokens rather than pixel-level predictions. This raises a critical question: How can we quantify the ability of VLMs to form spatial world modeling [17] in the presence of occlusion?
To address this, we introduce CAPTURe, Counting Amodally for Patterns Through Unseen REgions, a novel benchmark that tests a VLM’s world modeling and spatial reasoning abilities through the task of amodal counting, where models are prompted to count occluded objects by amodally completing a pattern. CAPTURe focuses on counting as it provides an objective and easy-to-verify output by comparing predicted counts with ground truth values. Moreover, patterned objects appear in various real-world domains, especially in man-made environments like parking lots, cities, and warehouses, where counting objects is often required. Fig. 1 illustrates the CAPTURe task. We show a VLM an image where objects are placed in a regular pattern (e.g., a 4x4 grid) with some objects occluded, and ask the model to count the total number of objects in the image assuming that the pattern continues behind the occlusion. The task requires handling occlusion, pattern recognition, and counting skills that exist in humans from a fairly young age [30, 27, 45], thus humans can easily answer such questions – indeed, we find that people can complete CAPTURe tasks with almost no error.
@@
By combining vision encoders with large language models (LLMs), VLMs have the potential to reason in a zero-shot way about visual inputs. To put this ability to the test and measure VLMs’ ability to reason about missing visual information, we evaluate four strong recent VLMs (GPT-4o, InternVL2, Molmo, and Qwen2VL) on CAPTURe. Our experiment results (Sec. 4) show that models generally struggle with the multiple aspects of the task, with high error rates on both CAPTURe and CAPTURe for occluded and unoccluded images. In contrast, we find that humans can perform the task easily: whereas model performance deteriorates as more objects in images are occluded, humans complete the task almost perfectly. We also compare VLMs to a vision-only model trained to count visible objects; while this model generally outperforms VLMs, its error is directly tied to the number of occluded objects – the more objects are occluded, the higher its error will be.
By objectively measuring VLMs’ spatial reasoning capabilities under occlusion, CAPTURe highlights an unexpected weakness in VLMs. We analyze this weakness by providing the model with additional clues and information. Specifically, we test to what degree the VLMs’ failure stems from an inability to integrate visual information by providing it with a text-based representation of the visible objects in the image in the form of object coordinates; here, VLMs perform substantially better, indicating that their poor performance on CAPTURe stems partly from an inability to count objects in images, rather than an inability to count more generally. Our findings align with previous work, which similarly finds that VLMs struggle to count in images [33, 22, 42]. We also test the degree to which VLM errors stem from an inability to form a world model by providing it with auxiliary information (the coordinates of the occluded objects in text, or inpainting the occluded regions). We find that VLMs perform substantially better with this auxiliary information, suggesting that VLMs are partly limited by their inability to imagine the missing visual information. Addressing these gaps is critical for VLMs to function effectively in real-world scenarios, where visual reasoning often involves occlusions – whether counting stadium seats, components on production lines, or buildings in neighborhoods. We hope that our work will foster future research on improving the world modeling capabilities of VLMs.
@@ 2.1 Task Overview
CAPTURe tests VLMs on occlusion reasoning, pattern recognition, and counting of both visible and occluded objects. VLMs already achieve high accuracy in classifying single, occluded objects [20]. Thus, we also argue that VLMs have the potential to perform well on CAPTURe’s challenging task because their proficiency in handling occlusion ought to enable them to recognize occluded objects and reason accordingly. All images in CAPTURe contain a pattern. This makes the task solvable for models and people – if the objects were not placed in a pattern, it would be unreasonable to expect models to infer the position of the occluded objects. For example, given an image of a random pile of coins with a region occluded, it is not easy to infer whether the occluded region contains no coins or contains roughly the same amount as the rest of the pile. For this task, the patterns considered are all regular and fairly small, e.g. grids, circles, triangles, and other regular shapes – see Fig. 2 for further examples. The last step of CAPTURe is counting, asking the model to provide an objectively measurable output. In addition to VLMs, we also test CountGD [3], a state-of-the-art object detection-based counting method, finding that it fails to account for the occluded scenario, as its training entails solely predicting the visible, unoccluded objects in the image.
@@
where represents the actual values, represents the predicted values, and is the number of observations. sMAPE is capped at 100%, providing a fixed range. This makes sMAPE ideal for challenging tasks like ours, as we can penalize responses that fail to produce an answer with a maximum error of 100%. For a justification of sMAPE over other metrics, see Sec. A.1.
@@ 2.2 Dataset
We introduce a set of real images with patterns to test amodal counting in naturalistic settings. The original images and annotations come from the FSC-147 dataset [37], a diverse counting dataset with manual annotations for the number of target objects and all object bounding boxes in each image. FSC-147 contains a diverse array of objects, with 6146 real-world images across 147 object categories. We filter FSC-147 for images that contain identifiable and regular patterns of objects and manually overlay a black box to occlude some objects, resulting in 924 images. Filtering is first performed with GPT-4o and then manually verified; we also manually verify that determining objects despite the occlusion is feasible. For each example, we maintain both occluded and unoccluded versions. Further details on CAPTURe can be found in Appendix B.
@@ 2.3 Statistics and Examples
Fig. 2 shows examples from CAPTURe and CAPTURe paired with their corresponding answers from GPT-4o and their ground truth answers. These examples show the range of objects and patterns in the dataset and highlight the task’s feasibility for humans. Tab. 1 reports summary statistics for CAPTURe, including the number of images and object types, as well as the mean number of occluded and total objects in both splits of CAPTURe. The number of objects in CAPTURe is shown in Fig. 3, where most images have between 0 and 30 objects. On CAPTURe, the maximum number of objects is 15, and CAPTURe images generally have 1-6 occluded objects (shown in Fig. 4, as further occlusion could make the count unresolvable).
@@ 3.1 Models
We experiment with GPT-4o [28], Intern-VL2-Llama3-8B [9, 10], Qwen2-VL-7B [41], MiniCPM-o 2.6 [47], and Kimi-VL-A3B [40] for their high scores on other VLM tasks [29]. We add Molmo 7B-D [13], because of its ability to “point and count,” giving it a potential advantage on CAPTURe. Specifically, Molmo is trained on millions of examples that directly ground text to 2D coordinates (or “points”) in images. This allows Molmo to directly point to image coordinates and count more easily by pointing to several objects. All the VLMs feature a different language backbone and vision encoder to provide broad coverage of model architectures. To evaluate models, we provide the model with the name of the specific object to be counted and the explicit instruction to count fully visible objects and objects behind the occluding box (in the occluded images). For each model, we test ten prompts on a validation set of 100 images, selecting the best prompt for each model in each dataset section (CAPTURe/CAPTURe) and for each environment (occluded/unoccluded). We provide the selected prompts in Appendix D.
@@ 3.2 Answer Generation and Extraction
Given the complex nature of CAPTURe, we allow models to generate open-ended responses and then subsequently extract answers. Further details (including the maximum number of tokens) can be found in Sec. A.2.
Answer extraction. Empirically, we found that constraining the output to a specific format for ease of analysis negatively impacted benchmark performance. Therefore, we instead prompt models to generate freely and extract the final output number using a separate answer extractor based on Llama 3.1 8B [1]. This answer extractor takes the output from the model as input and prompts it to extract a single number representing the final answer. The answer extractor also identifies if an output failed to converge on a singular number answer and assigns a label to these examples. We mark such incomplete/incoherent model generations as ‘skipped’ questions and when calculating the error later, these responses are assigned the worst possible sMAPE score (100%). The answer extractor outputs were manually verified on 1000 outputs, and the extractor was found to be 100% accurate.
Human and object detection baselines. We also report the performance of humans and a recent counting model (CountGD [3]) as baselines to establish a point of reference for model performance. To confirm that humans can perform the CAPTURe task, we provided 100 randomly selected occluded examples each from the CAPTURe and CAPTURe subsets to 3 undergraduate students with no prior knowledge of the task.
@@ 4.1 Main Results on CAPTURe
GPT-4o | 14.75 | 9.71 |
InternVL2 | 32.90 | 17.57 |
Molmo | 32.49 | 17.73 |
Qwen2VL | 29.33 | 11.74 |
Avg. of 4 VLMs | 27.37 | 14.19 |
@@
Models consistently struggle with counting and perform worse on occluded images. We run the VLMs on the occluded and unoccluded versions of CAPTURe to discern whether occlusion significantly impacts model performance. Tab. 2 shows that all models struggle with counting generally, performing poorly on both splits. Moreover, we see that every model performs better on the unoccluded images. On average, the models perform 6.28% worse in CAPTURe occluded images and 4.85% worse in CAPTURe occluded images (in terms of absolute sMAPE), indicating increased difficulty from a standard counting task. The best model for both splits, GPT-4o, has an error rate of 14.75% on CAPTURe and a lower error rate of 9.71% on CAPTURe. Across both the real and synthetic split, GPT-4o’s error increases with occlusion, by 1.41% on the real data and 3.81% on the synthetic split. Interestingly, despite its fine-tuning on counting tasks, Molmo exhibits a sizable error rate of 32.5% on CAPTURe occluded images. The high error rates of VLMs indicate limited capabilities in visual understanding under occlusions, pattern recognition, and counting. We further analyze the source of these errors with oracle experiments in Sec. 4.3.
@@
Humans complete the task with almost no error. Tab. 3, evaluated on a 100-example subset of each split, confirms that humans complete the task with ease despite occlusion, with an sMAPE of 3.79% on CAPTURe and 0.92% on CAPTURe. On the same subset of examples, models performed 7 times worse on CAPTURe and 14 times worse on CAPTURe than humans, underscoring the gap between VLMs and humans in this task.
@@
Object detection-based baseline outperforms VLMs. We attempt the task with a strong object detection-based model to highlight that a standard counting approach will experience a greater loss going from unoccluded to occluded environments, as it cannot capture any occluded objects, i.e. cannot reason. We choose CountGD [3], the top solution for unoccluded counting on FSC-147, on which it was trained. Because we draw our images from FSC-147’s train and test sets, and CountGD trains on FSC-147, we only evaluate CountGD on the subset of our data sourced from the FSC-147 test set, consisting of 149 images.
@@
We find that CountGD deteriorates by 7.19% on occluded images, increasing from 3.15% sMAPE to 10.34% as observed in Fig. 5. As expected, CountGD outperforms all VLMs on the unoccluded split as it is trained for counting on FSC-147. CountGD also outperforms the VLMs on the occluded split, reinforcing that only counting the visible objects is a hard-to-beat baseline. However, the drop in performance with occlusion is greater than the average VLM’s drop, highlighting a disadvantage of non-reasoning solutions on CAPTURe: their error is necessarily tied directly to the number of occluded objects and they cannot address the task on their own, whereas a VLM might be able to infer missing objects via reasoning.
@@
Hybrid VLM counting systems improve performance. Finding that CountGD is far better at counting visible objects than VLMs, we leverage the advantage that CountGD has by feeding its visible object count information to the VLMs as part of the prompt. As expected, Fig. 5 illustrates that there is a considerable decrease in error when CountGD and the VLMs are combined. However, this hybrid system still performs worse than CountGD alone, indicating VLMs are still subpar even at counting just occluded objects (as further reinforced by Sec. C.3).